2x-3y=12
x=4y+1
How do I solve the sytem of equation using the subsitution method? Please show the work!

Answer by christinaaa192(28)

first subsitute x=4y+1
2(4y+1)-3y=12 =
8y+2-3y=12 =
5y+2=12 =
5y=10 =
y=2
now that we know what y is (y=2), we can solve for x
2x-3(2)=12 =
2x-6=12
2x=18
x=9

i hope this helps